Find the zeroth order composite approximation for the boundary value problem ey + (cos x)y + (sin x)y=0 y(0) = 0 y ( ) =1.
Answer #1

for zeroth approximation, y = y t c for Thus, (Cosx) y + (Sinxly 20 .. outer solution ? - Cosxo so boundary layer will forsen

} Taylor expansion We know, Cosx = 1- och t Siurar - x² 3 cos Cen2 = 1+ OCC?) Sin Con = oce Put Y Y.C=0 =2 7+2=0 + d2 - - dn